# Quantitative Propagation of Chaos for 2D Viscous Vortex Model with General Circulations on the Whole Space

### v2

@@ Abstract.
We derive quantitative propagation of chaos in the sense of relative entropy for the 2D viscous vortex model with general circulations, approximating the vorticity formulation of the 2D Navier-Stokes equation on the whole Euclidean space, which is an extension of the previous works [17, 30]. We improve our previous results in [16] to the more general case that the vorticity may change sign, allowing the circulations to be in different orientations. We also extend the results of [55] to the whole space to obtain the sharp local propagation of chaos in the high viscosity regime. Both results of convergence rates are global-in-time on any finite time horizon and optimal among existing literature, thanks to the sharp logarithmic growth estimates and a new ODE hierarchy and iterated integral estimates.
Key words and phrases:
@@ 1. Introduction
The particles are labeled with the index $i$ ranging from 11 11 to $N$, with $X_{i}(t)\in{\mathbb{R}}^{2}$ and ${\mathcal{M}}_{i}/N\in{\mathbb{R}}$ representing, respectively, the position and circulation of the $i$-th vortex.
The displacement of the particles is determined by the combination effect of two factors. The first one is given by the pairwise interaction force with some weights given by the circulations of other particles. The different cases of circulations with ${\mathcal{M}}_{i}>0$ and ${\mathcal{M}}_{i}<0$ represent the two different orientations of the point vortices. We mention that the scale $1/N$ that appears in the vortex circulation is in the sense of mean-field scaling, keeping the whole circulation of the system bounded. The second one is given by the extra additive noise, driven by $N$ independent copies of the standard 2D Brownian motion, which are denoted by $\{B_{i}(t)\}_{1\leq i\leq N}$. The constant $\sigma>0$ represents the viscosity of the system dynamics and measures the intensity of the diffusion.
The interacting kernel $K$ has many different options arising from various natural physical models. In our study, this kernel is given by the famous 2D Biot-Savart law in fluid mechanics,
@@
where $x=(x_{1},x_{2})\in{\mathbb{R}}^{2}$. This point vortex model was initiated in the study of Helmholtz [25] and Kirchhoff [33]. It was Osada [43] who first showed that, although the kernel has some singularity of $O(1/|x|)$ at the origin, which makes the interaction between two particles large enough when they become sufficiently close to each other, the particles actually never collide in the evolution process with probability one. Therefore, the singularity of the system will never be reached, and the well-posedness of (1.1) is justified. This allows us to use the convention that $K(0)=0$, and thus add the terms with $i=j$ to the summation in (1.1).
Throughout this article, we make the common assumption that ${\mathcal{M}}_{i}$ is independent of time $t$, which means that the circulation of each point vortex, once determined at the initial time $t=0$, remains unchanged as the system evolves in time. We also assume that $\{{\mathcal{M}}_{i}\}_{1\leq i\leq N}$ are $N$ independent copies of some random variable ${\mathcal{M}}$ compactly supported in $[-A,A]$ for some $A>0$. Typical examples include the uniform distribution $U(-A,A)$.
@@
This new point vortex system (1.2) is endowed with scaled circulations ~$\tilde{{\mathcal{M}}}_{j}={\mathcal{M}}_{j}/A$, which are supported in $[-1,1]$. Therefore, we can always change our target system to satisfy the additional assumption $A\leq 1$, at the price that the viscosity constant $\sigma$ is also rescaled by the constant $A$, which is consistent with our assumptions on $\sigma$ in the main results.
Our assumption made on ${\mathcal{M}}_{i}$ corresponds to the practical statistical procedure that we put some i.i.d. circulations on each point vortex to expect the convergence of the vorticity as $N\rightarrow\infty$ in some Law of Large Numbers sense. This setting is a generalized version of the previous work of the authors [16], where the point vortex system has fixed circulations ${\mathcal{M}}_{i}\equiv 1$ at each vortex $X_{i}$ and thus the particle system is symmetric in position $X_{i}$. Note that this is not the case in this article, since now we can only assume symmetry on the full information $Z_{i}=({\mathcal{M}}_{i},X_{i})$ for each particle.
We are interested in the mean-field limit dynamics of the point vortex system (1.1) when $N\rightarrow\infty$ in some appropriate sense of statistical mechanics. It is well-known that the limit self-consistent McKean-Vlasov system writes
@@
[…]
@@
The relation between the particle system (1.1) and the limit McKean-Vlasov system (1.3) as $N\rightarrow\infty$ lies in the regime of the mean-field theory and is described by the so-called propagation of chaos argument. For our stochastic point vortex system approximating the limit 2D Navier-Stokes equation, the first such result dates back to Osada [43], under the assumption of large viscosity. Propagation of chaos in the $L^{1}$ sense with positive viscosity was first established in the seminal paper of Fournier-Hauray-Mischler [17], where they studied the entropy and Fisher information of the point vortex system and used a compactness argument, but without any explicit convergence rate. This article is devoted to quantitative propagation of chaos for the point vortex system (1.1). For detailed information and comprehensive results of recent developments in mean-field theory and propagation of chaos, we refer to reviews such as [53, 27, 18, 29].
@@ 1.1. Liouville Equation, BBGKY Hierarchy and Limit Equation
which corresponds to the choice of i.i.d. initial data of the point vortex system (1.1): $\{\mathcal{M}_{i},X_{i}(0)\}_{1\leq i\leq N}$ are $N$ independent copies of $(\mathcal{M},X_{0})$ with law $g_{0}$. We also define the $k$-marginal of the joint law of the particle system by
@@
We can relax the factorization assumption on the initial data $G_{0}^{N}$. For our purposes below, it is sufficient to assume that the marginals of all orders of the initial distribution $G_{0}^{N,k}$ are close to their factorized counterparts $g_{0}^{\otimes k}$ in the sense of relative entropy as below:
$H(G_{0}^{N,k}|g_{0}^{\otimes k})=\int_{\mathbb{D}^{k}}G_{0}^{k}\log\frac{G_{0}^{k}}{g_{0}^{\otimes k}}\,\mathrm{d}Z^{k}\lesssim\frac{k^{2}}{N^{2}}.$
We choose the simplest factorization assumption here, which corresponds to that the relative entropy $H(G_{0}^{N,k}|g_{0}^{\otimes k})$ vanishes, but our results can be easily applied to the above more general case.
Applying Itô formula to the particle system (1.1), one concludes that the joint law $G_{t}^{N}$ solves the following Liouville equation:
@@
As for the limit system (1.3), by applying Itô formula, we also obtain that the limit density $g_{t}(m,x)$ solves the following nonlinear Fokker-Planck equation
@@
Finally we define our normalized relative entropy between the joint distribution of the particle system and the $k$-factorized law of the limit system, which will be the major quantity that we shall estimate.
Definition 1.2 (Normalized Relative Entropy).
@@ 1.2. Main Results
For the global case, inspired by the groundbreaking paper [30], we consider the global normalized relative entropy between the joint law of the particle system and the factorized law of the limit system, which controls the $L^{1}$ distance of marginals between two systems by the Csiszár-Kullback-Pinsker (CKP) inequality. By estimating the time evolution of the relative entropy, using the Law of Large Numbers and Large Deviation type results introduced by [30], we are able to close the Grönwall loop and arrive at a global optimal decay rate, hence proving the strong propagation of chaos result. We mention that in this case we prove our results for all positive viscosity $\sigma>0$. This part of our main results is an extension of the previous work [17] for qualitative propagation of chaos and the previous work by the authors [16] for the special case ${\mathcal{M}}_{j}\equiv 1$ and by Shao-Zhao [52] for the torus case.
@@
$\|\omega^{N,k}-\omega^{\otimes k}\|_{TV}\leq\|G^{N,k}-g^{\otimes k}\|_{TV}\leq\frac{\sqrt{k}}{\sqrt{N}}M(1+t)^{M}$
@@
to derive the local estimates from the global ones. Namely, this will provide the convergence rate in total variation as
$\|\omega^{N,k}-\omega^{\otimes k}\|_{TV}\lesssim\frac{\sqrt{k}}{\sqrt{N}},$
@@
$\|\omega^{N,k}-\omega^{\otimes k}\|_{TV}\leq\frac{k}{N}M(1+t)^{M}$
@@ 1.3. Related Works
Propagation of chaos and convergence to mean-field limit for large systems of interacting particles, widely appearing in the analysis of PDEs and probability theory, have been extensively studied over the last decades, both for the first-order and second-order systems. The derivation of some effective equations to describe the average or statistical behavior of large interacting particle systems can be traced back to Maxwell [40] and Boltzmann [2]. Kac [31] first raised the basic notions of chaos and proved some partial propagation of chaos results for the simplified model of the Boltzmann equation in dimension 11 11. Related studies have been far beyond the range of kinetic theory. The fundamental mathematical setting of this problem was set up and basically resolved by McKean [41] and Dobrushin [14] using the classical coupling method. Since then people began to pay attention to the much more difficult models, say systems with singular interacting kernels, instead of the regular (bounded Lipschitz) kernels. As for our point vortex model approximating the famous 2D Navier-Stokes/Euler equation, the first result dates back to Osada [43] with bounded initial data and large viscosity. Later Fournier-Hauray-Mischler [17] extended the result to all positive viscosity cases, while the initial data is only assumed to have finite moments and entropy, using the compactness argument and the entropy estimate.
@@
The modulated energy method introduced by Serfaty [50] is another effective tool for resolving propagation of chaos for the singular kernel case, especially the Coulomb/Riesz cases. By constructing a Coulomb/Riesz based metric between the empirical measure of the $N$-particle system and the limit density. Serfaty [50] established the quantitative propagation of chaos for the first-order Coulomb and super-Coulomb cases without noise in the whole space, in the sense of controlling the weak convergence rate when acting on test functions. This method has been extended to various problems, including Rosenzweig [45] for 2D incompressible Euler equation, Nguyen-Rosenzweig-Serfaty [42] for Riesz-type flows, Rosenzweig-Serfaty [46] for uniform-in-time estimates for sub-Coulomb cases, and De Courcel-Rosenzweig-Serfaty [13] for periodic Riesz cases. The key estimates appearing in the control of modulated energy, say the commutator type estimates, have been thoroughly examined in the recent work Rosezweig-Serfaty [49]. The combination of modulated energy and relative entropy leads to the modulated free energy method, which is applied to the Coulomb/Riesz kernels on the torus and more interestingly the first time to the attractive kernel cases by Bresch-Jabin-Wang [6, 5, 7] for the 2D log gas and the 2D Patlak-Keller-Segel model. See also the recent work by De Courcel-Rosenzweig-Serfaty [12] for further results on uniform-in-time propagation of chaos and phase transition for attractive log gas. In the recent work Cai-Feng-Gong-Wang [9], the authors and coauthors are also able to apply the modulated free energy method to resolve the 2D Coulomb case on the whole space. Some generation of chaos results are also obtained partially by Rosenzweig-Serfaty [47, 48] using the time-uniform LSI for the limit density or the modulated Gibbs measure. We refer to the recent lecture notes by Serfaty [51] for more complete discussions on the Coulomb/Riesz gases concerning above topics.

### v3

@@ Abstract.
We derive quantitative propagation of chaos in the sense of relative entropy for the 2D viscous vortex model with general circulations, approximating the vorticity formulation of the 2D Navier-Stokes equation on the whole Euclidean space. Our results work on the general setting that the vortices are positioned on the whole space ${\mathbb{R}}^{2}$ and that the circulations are allowed to be in different magnitudes and orientations, which can be adapted to general unconfined realistic fluids with vorticity that may change sign. We provide explicit convergence rates which are optimal in $N$ and optimal in $t$ among existing literature. The key technical tools, which are our major novelty, are the sharp logarithmic growth estimates and a new ODE hierarchy and iterated integral estimates.
@@ 1. Introduction
The particles are labeled with the index $i$ ranging from $1$ to $N$, with $X_{i}(t)\in{\mathbb{R}}^{2}$ and ${\mathcal{M}}_{i}/N\in{\mathbb{R}}$ representing, respectively, the position and circulation of the $i$-th vortex.
The circulation ${\mathcal{M}}_{i}/N$ is an important quantity in the point vortex model. It measures the strength of the $i$-th vortex. The different cases of circulations with ${\mathcal{M}}_{i}>0$ and ${\mathcal{M}}_{i}<0$ represent the two different orientations of the point vortices. It is a very common physical phenomenon in real-world flows to have vortices of different signs. A typical example is a vortex dipole, a pair of closely spaced vortices with circulations of equal magnitude but opposite sign. We mention that the scale $1/N$ that appears in the vortex circulation is in the sense of mean-field scaling, keeping the whole circulation of the system bounded.
The displacement of the particles is determined by the combination effect of two factors. The first one is given by the pairwise interaction force with some weights given by the circulations of other particles. The interacting kernel $K$ has many different options arising from various natural physical models. In our study, this kernel is given by the famous 2D Biot-Savart law in fluid mechanics,
@@
where $x=(x_{1},x_{2})\in{\mathbb{R}}^{2}$. The second one is given by the extra additive noise, driven by $N$ independent copies of the standard 2D Brownian motion, which are denoted by $\{B_{i}(t)\}_{1\leq i\leq N}$. The constant $\sigma>0$ represents the viscosity of the system dynamics and measures the intensity of the diffusion.
The point vortex model (1.1) has significant physical importance and meanings, especially in fluid mechanics and statistical mechanics. In fluid mechanics, the motion of a viscous incompressible fluid is classically described by the Navier-Stokes equation. The notion of vorticity and the corresponding vorticity formulation of the 2D Navier-Stoke equation, which describe the local rotation of the velocity field, are powerful and intuitive to understand complex fluid flows in meteorology or oceanography, for instance. To better understand the vorticity function, a long-standing and fruitful approach in mathematical physics and applied mathematics is to approximate the vorticity equation by a discrete interacting point vortex system. The discretized version of the vorticity is useful and effective in both physically modeling and numerically simulating the 2D Navier-Stokes equation.
Such point vortex model was initiated in the study of Helmholtz [25] and Kirchhoff [33]. Mathematically, it was Osada [43] who first showed that, although the kernel has some singularity of $O(1/|x|)$ at the origin, which makes the interaction between two particles large enough when they become sufficiently close to each other, the particles actually never collide in the evolution process with probability one. Therefore, the singularity of the system will never be reached, and the well-posedness of (1.1) is justified. This allows us to use the convention that $K(0)=0$, and thus add the terms with $i=j$ to the summation in (1.1).
Throughout this article, we make the common assumption that ${\mathcal{M}}_{i}$ is independent of time $t$, which means that the circulation of each point vortex, once determined at the initial time $t=0$, remains unchanged as the system evolves in time. We also assume that $\{{\mathcal{M}}_{i}\}_{1\leq i\leq N}$ are $N$ independent copies of some random variable ${\mathcal{M}}$ compactly supported in $[-A,A]$ for some $A>0$. Typical examples include the uniform distribution $U(-A,A)$. Our assumption made on ${\mathcal{M}}_{i}$ corresponds to the practical statistical procedure that we put some i.i.d. circulations on each point vortex to expect the convergence of the vorticity as $N\rightarrow\infty$ in some Law of Large Numbers sense.
@@
This new point vortex system (1.2) is endowed with scaled circulations $\tilde{{\mathcal{M}}}_{j}={\mathcal{M}}_{j}/A$, which are supported in $[-1,1]$. Therefore, we can always change our target system to satisfy the additional assumption $A\leq 1$, at the price that the viscosity constant $\sigma$ is also rescaled by the constant $A$, which is consistent with our assumptions on $\sigma$ in the main results.
We are interested in the mean-field limit dynamics of the point vortex system (1.1) when $N\rightarrow\infty$ in some appropriate sense of statistical mechanics. It is well-known in the mean-field theory that the limit self-consistent McKean-Vlasov system writes
@@
represents the joint density of ${\mathcal{M}}$ and $X_{t}$. By the Itô formula, it is straightforward to check that the vorticity function $\omega_{t}$ satisfies the famous vorticity formulation of the 2D Navier-Stokes equation in fluid mechanics:
@@
Therefore, once we rigorously justify the mean-field convergence from the particle system (1.1) to the limit McKean-Vlasov system (1.3) or its corresponding Fokker-Planck equation (1.4) in some appropriate sense, then we can show that the point vortex system (1.1) is effective in mathematically and physically modeling or numerically simulating the 2D Navier-Stokes equation. Moreover, if we can quantify the mean-field convergence rate or establish the convergence in a better sense, the point vortex system is shown more powerful. Also, the mean-field convergence is also a key tool for understanding the large point vortex system when $N$ becomes extremely large. In such cases, it is difficult to track and analyze the trajectory and the behavior of each particle. Therefore, we turn to the statistical behavior of the particles, which can be well approximated by the limit density $\omega_{t}$. The arguments above also apply to more general models, including many stochastic interacting particle systems in higher dimensions.
The relation between the particle system (1.1) and the limit McKean-Vlasov system (1.3) as $N\rightarrow\infty$ lies in the regime of the mean-field theory. It is expected to justify that, if the initial empirical measure $\frac{1}{N}\sum_{i=1}^{N}{\mathcal{M}}_{i}\delta_{X_{i}(0)}$ converges weakly to $\omega_{0}$, then for any $t>0$, the empirical measure $\frac{1}{N}\sum_{i=1}^{N}{\mathcal{M}}_{i}\delta_{X_{i}(t)}$ of the particle system (1.1) also converges weakly to $\omega_{t}$, where $\omega_{t}$ solves the 2D Navier-Stokes equation (1.4). In the language of Kac’s molecular chaos [31], this is called propagation of chaos. An equivalent argument can be stated in the PDE context. For any fixed $k$, denote by $G_{t}^{N,k}$ the probability density function of $({\mathcal{M}}_{1},X_{1},\cdots,{\mathcal{M}}_{k},X_{k})$ at time $t$ and let $\omega_{t}^{N,k}={\mathbb{E}}_{{\mathcal{M}}_{1}\cdots{\mathcal{M}}_{k}}G_{t}^{N,k}$ be the $k$-particle vorticity function. We expect that if $\omega_{0}^{N,k}$ converges weakly to $\omega_{0}^{\otimes k}$ initially, then for any $t>0$, $\omega_{t}^{N,k}$ also converges weakly to $\omega_{t}^{\otimes k}$. This argument shows that, although the particles cannot be independent when $t>0$ due to the pairwise interaction, any given fixed $k$ particles are asymptotically independent as $N\to\infty$, recovering the independence in some sense. Moreover, this convergence can be quantified using the distances between probability density functions, which is referred to quantitative propagation of chaos. For detailed information and comprehensive results of recent developments in mean-field theory and propagation of chaos, we refer to reviews such as [53, 27, 18, 29].
For our stochastic point vortex system approximating the limit 2D Navier-Stokes equation, the first such result dates back to Osada [43], under the assumption of large viscosity. Propagation of chaos in the $L^{1}$ sense with positive viscosity was first established in the seminal paper of Fournier-Hauray-Mischler [17], where they studied the entropy and Fisher information of the point vortex system and used a compactness argument, but without any explicit convergence rate. This article is devoted to quantitative propagation of chaos for the point vortex system (1.1).
@@ 1.1. Liouville Equation, BBGKY Hierarchy and Limit Equation
which corresponds to the choice of i.i.d. initial data of the point vortex system (1.1): $\{\mathcal{M}_{i},X_{i}(0)\}_{1\leq i\leq N}$ are $N$ independent copies of $(\mathcal{M},X_{0})$ with law $g_{0}$. Under such initial distribution, the particles in the point vortex system (1.1) are initially well-prepared or well-distributed with probability one, therefore creating no collisions when the system evolves, as proved by Osada [43]. We also define the $k$-marginal of the joint law of the particle system by
@@
We define the normalized relative entropy between the joint distribution of the particle system and the $k$-factorized law of the limit system, which will be the major quantity that we shall estimate.
Definition 1.1 (Normalized Relative Entropy).
@@
We can relax the factorization assumption on the initial data $G_{0}^{N}$. For our purposes below, it is sufficient to assume that the marginals of all orders of the initial distribution $G_{0}^{N,k}$ are close to their factorized counterparts $g_{0}^{\otimes k}$ in the sense of the normalized relative entropy in Definition 1.1 as below:
$H_{k}(G_{0}^{N,k}|g_{0}^{\otimes k})=\frac{1}{k}\int_{\mathbb{D}^{k}}G_{0}^{k}\log\frac{G_{0}^{k}}{g_{0}^{\otimes k}}\,\mathrm{d}Z^{k}\lesssim\frac{k}{N^{2}}.$
We choose the simplest factorization assumption here, which corresponds to that the normalized relative entropy $H_{k}(G_{0}^{N,k}|g_{0}^{\otimes k})$ vanishes, but our results can be easily applied to the above more general case.
Applying the Itô formula to the particle system (1.1), one concludes that the joint law $G_{t}^{N}$ solves the following Liouville equation:
@@
As for the limit system (1.3), by applying the Itô formula, we also obtain that the limit density $g_{t}(m,x)$ solves the following nonlinear Fokker-Planck equation
@@ 1.2. Main Results
For the global case, inspired by the groundbreaking paper [30], we consider the global normalized relative entropy between the joint law of the particle system and the factorized law of the limit system, which controls the $L^{1}$ distance of marginals between two systems by the Csiszár-Kullback-Pinsker (CKP) inequality. By estimating the time evolution of the relative entropy, using the Law of Large Numbers and Large Deviation type results introduced by [30], we are able to close the Grönwall loop and arrive at a global optimal decay rate, hence proving the strong propagation of chaos result. We mention that in this case we prove our results for all positive viscosity $\sigma>0$.
@@
$\|\omega^{N,k}-\omega^{\otimes k}\|_{L^{1}({\mathbb{R}}^{2k})}\leq\|G^{N,k}-g^{\otimes k}\|_{L^{1}(\mathbb{D}^{k})}\leq\frac{\sqrt{k}}{\sqrt{N}}M(1+t)^{M}$
@@
to derive the local estimates from the global ones. Namely, this will provide the convergence rate in $L^{1}$ distance as
$\|\omega^{N,k}-\omega^{\otimes k}\|_{L^{1}({\mathbb{R}}^{2k})}\lesssim\frac{\sqrt{k}}{\sqrt{N}},$
@@
$\|\omega^{N,k}-\omega^{\otimes k}\|_{L^{1}({\mathbb{R}}^{2k})}\leq\frac{k}{N}M(1+t)^{M}$
@@
1.3. Discussion on the Model
In this subsection, we gather some discussions on the vortex model (1.1) in our main results, concerning the main motivation to include the general circulations and the whole space as a state space and also the possible limitations on bounded domains. We emphasize that these conditions are motivated both physically and mathematically.
The appearance of general circulations ${\mathcal{M}}_{i}/N$ essentially improves the universality of the vortex model, making it capable of approximating more general fluid flows. In the realistic setting, vortex models such as a dipole naturally contain vortices of different signs. Also in the real-world flows, it is natural to expect that the strength and the direction of the local rotation in the fluid are not constant everywhere. Therefore, such cases require a point vortex model with different circulations for a suitable approximation. Mathematically, as explained in Remark 1.2, the simpler model with fixed circulations ${\mathcal{M}}_{i}\equiv 1$, although being easier for computation, requires that the vorticity $\omega_{t}$ is given by a probability density function , or at least a fixed-sign function. This limitation prevents the approximation of physically more general and realistic vorticity functions.
We have also chosen the whole space ${\mathbb{R}}^{2}$ as a state space for the point vortex model, instead of a compact manifold like the torus on which many previous studies worked. Namely, we assume on the position of the vortices $X_{i}\in{\mathbb{R}}^{2}$ rather than $X_{i}\in\mathbb{T}^{2}$. The physical consideration is also to restore the generality of our model. Many real fluids, like the ocean or a large body of water, exists in extremely large domains. It is somehow artificial and unrealistic to constrain them in a compact manifold. Moreover, confining the fluid and the model to a torus imposes an extra periodicity which is non-physical, and also prevents us from studying how the vorticity spreads and decays at infinity. It also requires new mathematical techniques to deal with the vorticity function on the whole space, namely the sharp logarithmic growth estimates in our results, in order to handle the quantities like $\nabla\log g$ which are typically unbounded on the whole space but are easier to estimate on the torus. This provides new insights for the study of the singular 2D Biot-Savart kernel and the similar behavior of the fundamental solution with the well-known Oseen vortices.
There are possible limitations when we work on the realistic bounded domains. On a bounded domain, the Navier-Stokes equation requires physical boundary conditions, most commonly the no-slip condition (the fluid velocity is zero at the boundary). But the vorticity formulation does not behave well with these boundary conditions. In a real viscous fluid, the no-slip condition at the boundary provides a primary mechanism for vorticity generation. Vorticity is created at the boundary and diffuses into the flow. Moreover, the boundary condition may also be not compatible with the mean-field scaling. The entire derivation of the mean-field limit relies on the fact that the interaction force on a given vortex is a sum of $N-1$ terms, each of order $1/N$. On a bounded domain, however, the interaction with the boundary does not necessarily respect this scaling. The influence of the boundary may become a dominant non-mean-field effect.
@@ 1.4. Related Works
Propagation of chaos and convergence to mean-field limit for large systems of interacting particles, widely appearing in the analysis of PDEs and probability theory, have been extensively studied over the last decades, both for the first-order and second-order systems. The derivation of some effective equations to describe the average or statistical behavior of large interacting particle systems can be traced back to Maxwell [40] and Boltzmann [2]. Kac [31] first raised the basic notions of chaos and proved some partial propagation of chaos results for the simplified model of the Boltzmann equation in dimension $1$. Related studies have been far beyond the range of kinetic theory. The fundamental mathematical setting of this problem was set up and basically resolved by McKean [41] and Dobrushin [14] using the classical coupling method. Since then people began to pay attention to the much more difficult models, say systems with singular interacting kernels, instead of the regular (bounded Lipschitz) kernels. As for our point vortex model approximating the famous 2D Navier-Stokes/Euler equation, the first result dates back to Osada [43] with bounded initial data and large viscosity. Later Fournier-Hauray-Mischler [17] extended the result to all positive viscosity cases, while the initial data is only assumed to have finite moments and entropy, using the compactness argument and the entropy estimate.
@@
The modulated energy method introduced by Serfaty [50] is another effective tool for resolving propagation of chaos for the singular kernel case, especially the Coulomb/Riesz cases. By constructing a Coulomb/Riesz based metric between the empirical measure of the $N$-particle system and the limit density. Serfaty [50] established the quantitative propagation of chaos for the first-order Coulomb and super-Coulomb cases without noise in the whole space, in the sense of controlling the weak convergence rate when acting on test functions. This method has been extended to various problems, including Rosenzweig [45] for 2D incompressible Euler equation, Nguyen-Rosenzweig-Serfaty [42] for Riesz-type flows, Rosenzweig-Serfaty [46] for uniform-in-time estimates for sub-Coulomb cases, and Chodron De Courcel-Rosenzweig-Serfaty [12] for periodic Riesz cases. The key estimates appearing in the control of modulated energy, say the commutator type estimates, have been thoroughly examined in the recent work Rosezweig-Serfaty [49]. The combination of modulated energy and relative entropy leads to the modulated free energy method, which is applied to the Coulomb/Riesz kernels on the torus and more interestingly the first time to the attractive kernel cases by Bresch-Jabin-Wang [6, 5, 7] for the 2D log gas and the 2D Patlak-Keller-Segel model. See also the recent work by Chodron De Courcel-Rosenzweig-Serfaty [13] for further results on uniform-in-time propagation of chaos and phase transition for attractive log gas. In the recent work Cai-Feng-Gong-Wang [9], the authors and coauthors are also able to apply the modulated free energy method to resolve the 2D Coulomb case on the whole space. Some generation of chaos results are also obtained partially by Rosenzweig-Serfaty [48, 47] using the time-uniform LSI for the limit density or the modulated Gibbs measure. We refer to the recent lecture notes by Serfaty [51] for more complete discussions on the Coulomb/Riesz gases concerning above topics.